OpenStudy (anonymous):

How would I put -y + 5 = -2x into slope-int form?

OpenStudy (anonymous):

\[ -y+5=-2x\\ -y=-2x-5\\ y=2x+5 \]

OpenStudy (anonymous):

put the equation in the formula y=mx+b

OpenStudy (anonymous):

why does the negative 2 change to a positive?

OpenStudy (anonymous):

@chungerforever multiply through by -1 to make y positive :-)

OpenStudy (anonymous):

because you distribute it

OpenStudy (anonymous):

okay, and if i wanted to graph it would I start at 5 then rise 2 and over 1?

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

thank you! (: